# Forgetting Order of Continual Learning: Examples That are Learned First are Forgotten Last

### v1

@@ Abstract
Catastrophic forgetting poses a significant challenge in continual learning, where models often forget previous tasks when trained on new data. Our empirical analysis reveals a strong correlation between catastrophic forgetting and the learning speed of examples: examples learned early are rarely forgotten, while those learned later are more susceptible to forgetting. We demonstrate that replay-based continual learning methods can leverage this phenomenon by focusing on mid-learned examples for rehearsal. We introduce Goldilocks, a novel replay buffer sampling method that filters out examples learned too quickly or too slowly, keeping those learned at an intermediate speed. Goldilocks improves existing continual learning algorithms, leading to state-of-the-art performance across several image classification tasks.
@@ 1 Introduction
The surge in deploying deep models in real-world scenarios has heightened the need for models capable of adjusting to changing environments. Continual learning (see extensive reviews, e.g [10, 19, 37]), which involves training deep learning models on new data even after mastering previous tasks, is crucial in this context. However, continual learning encounters a significant obstacle called catastrophic forgetting [12, 24, 32]: as models learn from new data, they tend to forget previous tasks, resulting in decreased performance on those tasks. Despite extensive research efforts, catastrophic forgetting persists as a central challenge in continual learning.
We begin by empirically examining catastrophic forgetting, identifying which samples deep models tend to forget when trained on new tasks. Our findings reveal a last-in-first-out pattern, where examples learned later are forgotten more readily than those learned earlier. Fig. 1 visualizes this by showing which examples are likely to be forgotten or remembered over time. This phenomenon is connected to the simplicity bias [48, 50] of neural networks, where networks learn simple examples faster than complex ones. However, complex examples are more likely to be forgotten, while simple examples are more likely to be remembered.
This observation has an important consequence: by measuring how fast examples are being learned by the model, we can predict in advance which examples are more susceptible to forgetting. This holds true regardless of the distribution of the new task introduced to the model. Furthermore, we demonstrate that models showing varying degrees of forgetting do so consistently: models with less catastrophic forgetting recall most of the examples retained by those with more forgetting, with the additional ability to remember slightly more complex examples.
Replay-based continual learning methods use a memory buffer to address catastrophic forgetting, storing a small subset of original task examples. Typically, this buffer is uniformly sampled from previous tasks, and different continual learning techniques vary in how they utilize it. Our findings indicate that while a replay buffer helps mitigate forgetting, a similar pattern emerges: examples learned later are forgotten sooner. Increasing buffer size or improving the underlying model’s architecture can alleviate more of the forgetting to some extent, but the fundamental phenomenon persists. This observation implies that when strengthening a continually trained model, we can predict in advance which examples it will better remember compared to its weaker version.
This raises the question: can one exploit this phenomenon and reduce forgetting by cleverly selecting samples for the replay buffer? Our observations reveal that including in the replay buffer more examples that the model can potentially remember given more practice can improve a large range of continual learning methods. Given the fixed buffer size, these examples replace the fastest-to-learn examples, which the model is likely to remember anyway, and the slowest-to-learn examples, which it is likely to forget. Leveraging these insights, we introduce a practical sampling method named Goldilocks. In constructing the replay buffer, Goldilocks eliminates examples learned too quickly or too slowly, ensuring the buffer comprises examples learned at just the right speed.
Integrating Goldilocks sampling into different competitive continual learning algorithms is computationally negligible, and results in a significant improvement across a diverse range of image classification tasks, surpassing state-of-the-art performance in many cases. Unlike prior replay-buffer sampling methods, which often exhibit effectiveness only within specific contexts or with particular algorithms, Goldilocks consistently performs well across all evaluated settings.
@@
Figure 1: Test examples that networks tend to either remember or forget. Networks were trained on 2 tasks from CIFAR-100. Examples from the first task that were learned quickest (left) and slowest (right) are shown. Slowest examples were often forgotten, while the quickest were seldom forgotten.
@@ 1.1 Our contribution
  1. 1.
    We demonstrate consistent catastrophic forgetting across diverse continual learning approaches, highlighting certain examples’ susceptibility to forgetting regardless of the specific continual learning framework.
  2. 2.
    Networks are more likely to forget examples learned during later stages of the learning than examples learned early on. This observation allows us to predict during training the likelihood a given example will be forgotten.
  3. 3.
    We introduce Goldilocks, a novel and efficient replay buffer sampling method. Goldilocks selectively samples examples learned midway through training, discarding examples that are too simple or too complex for the model. Goldilocks reduces catastrophic forgetting and improves the performance of a large variety of continual learning methods.
@@ 1.2 Related work
Existing studies primarily focus on mitigating catastrophic forgetting [25, 28, 29, 45, 47], rather than its characterization. Some works have explored catastrophic forgetting from a model perspective: [36, 42] showed differences in forgetting patterns across model layers, while [34, 40] demonstrated the impact of training hyper-parameters. [35] identified task complexity as a factor influencing forgetting rates. Our study differs by examining the relationship between learning speed and forgettability of examples, emphasizing a data-centric perspective.
Simplicity bias. Empirical evidence supports the hypothesis that neural networks initially learn simple models, gaining complexity over time [7, 14, 15, 20, 21, 22, 23, 38, 49, 54]. This phenomenon, known as simplicity bias [11, 48], is extended in our work to catastrophic forgetting, illustrating a reverse simplicity bias where complex examples are forgotten before simpler ones.
Forgetting dynamics. Several studies have directly explored the dynamics of forgetting. [31, 53] revealed that, even in non-continual settings, certain examples tend to be less forgettable during training. [33] demonstrated that the forgetting dynamics of noisy examples differ from those of clean labels and proposed adjusting the rehearsal process accordingly. Our results align with these findings, further delineating which examples are more prone to forgetting and retention, establishing connections between catastrophic forgetting and simplicity bias, and proposing a method to identify examples likely to be forgotten proactively. Additionally, our work illustrates how to leverage these insights to enhance the sampling of the replay buffer for various continual learning approaches.
Replay buffer sampling functions. Replay-based methods mitigate catastrophic forgetting by storing examples from previous tasks in a memory buffer of fixed size. Despite the availability of various sampling strategies [1, 3, 6, 44, 52, 55], their effectiveness often depends on specific conditions or algorithms. As a result, many replay-based methods resort to uniform sampling [5, 16, 25, 30, 41, 43, 46] in practice. Unlike previous methods, Goldilocks can improve replay-based methods across diverse tasks and scenarios.
2 Empirical analysis of catastrophic forgetting
@@
In this section, we empirically study catastrophic forgetting from a data-centric perspective, drawing a connection between the speed at which a model learns examples, and its likelihood to forget them.
@@ 2.1 Definitions
We examine the continual learning classification setting, where the training dataset $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ consists of examples $x\in\mathcal{X}$ and their corresponding labels $y\in\mathcal{Y}$. This dataset is divided into $T>1$ tasks, with each task $t\in[T]$ represented as $\mathcal{D}_{t}=\{\mathcal{X}_{t},\mathcal{Y}_{t}\}$, and $\mathcal{D}=\bigcup_{t=1}^{T}{D_{t}}$. There is no overlap between the data of different tasks. A deep model $f:\mathcal{X}\rightarrow\mathcal{Y}$ is sequentially trained on each task $\mathcal{D}_{t}$ from $t=1$ to $t=T$, with $E\in\mathbb{N}$ epochs per task. Each task also has a separate test set from the same distribution. During training on a task $\mathcal{D}_{t}$, the model $f$ is provided with a replay buffer $B$ of fixed-memory size, containing examples from previous tasks ${1,...,t-1}$. Notably, $|B|\ll|\mathcal{D}|$, meaning only a fraction of the data can be stored in the replay buffer. It’s worth mentioning that the buffer may be empty ($B=\emptyset$), representing training $f$ without a memory buffer.
Learning speed. Neural networks demonstrate a trend in learning examples, where certain examples are consistently learned before others across various deep neural models [2, 9, 17]. This orderliness is often connected to the simplicity bias phenomenon [18], as examples learned earlier tend to be simpler than those learned later. Rather than measuring the example’s complexity, which can be defined in many different ways, we use the speed at which examples are learned by the deep model, which can be measured efficiently and directly. Given a model $f$ and an example $(x,y)$ either in the train or test data, we define the learning speed of the example $(x,y)$ to be:
@@
Here, $f_{t}^{e}$ denotes the intermediate model $f$ after training for $e\in[E]$ epochs on task $\mathcal{D}_{t}$. This definition is based on the accessibility score introduced in [17], adapted to be computed by a single model rather than an ensemble, enabling efficient computation throughout the training of $f$. Intuitively, this score indicates how rapidly the model learns an example: an example correctly classified from the early stages of learning will sustain $f_{t}^{e}(x)=y$ for more epochs, resulting in a higher learning speed.
Computing the learning speed. To calculate the learning speed of each example, we maintain a boolean matrix, called the epoch-wise classification matrix, $M\in\left\{0,1\right\}^{E\times|\mathcal{D}_{t}|}$ throughout $f$’s training on task $\mathcal{D}_{t}$, indicating whether the model correctly classified each example during learning. Specifically, for example $(x_{i},y_{i})\in\mathcal{D}_{t}$, we have $M_{e,i}=\mathds{1}[f_{t}^{e}(x_{i})=y_{i}]$. An example’s learning speed is then the average of the $M$ across epochs. This matrix can be computed during the forward pass of the network, incurring minimal to no additional computation overhead. Furthermore, memory usage can be further reduced by directly computing the matrix’s mean during training.
@@ 2.2 Preliminaries
Class incremental vs. task incremental learning. While similar, task incremental learning distinguishes itself from class incremental learning by providing the task identity for each example during testing, enabling the use of the specific classifier of the classes within the task. As a result, performance in class incremental learning is typically much lower. Due to space constraints, our main paper exclusively presents results for task incremental learning. However, findings for class incremental learning, which align qualitatively with the task incremental case, are available in App. A.
Datasets. We investigated various image continual learning classification tasks using split versions of several image datasets, including CIFAR-10, CIFAR-100 [26], and TinyImagenet [27]. The data is split into $T$ tasks by partitioning the classes into $T$ equal-sized subsets. This partitioning is denoted as dataset-T. For example, splitting CIFAR-10 into 5 classes is denoted as CIFAR-10-5, comprising 5 tasks, each with 2 distinct classes. Unless otherwise specified, classes are divided into tasks according to the original order they appeared in the dataset (often alphabetically).
Architectures and training hyper-parameters. In our experiments without a replay buffer, unless stated otherwise, we trained ResNet-18 for $E=100$ epochs per task. We employed a base learning rate of $0.1$ with a cosine scheduler, SGD optimizer, momentum of $0.9$, and weight decay of $0.0005$. All networks were trained on NVIDIA TITAN X. These hyperparameters were chosen arbitrarily based on their performance during joint dataset training and were consistent across all experiments. We anticipate consistent qualitative results despite potential variations in these hyperparameters. When introducing a replay buffer, we employed an experience replay strategy [46], alternating batches of data from the new task and the replay buffer. When comparing against other replay buffer sampling methods, we maintained the same ResNet-18 settings. When integrating with other continual learning algorithms, we evaluated all methods within the framework of [4, 5], using the best hyperparameters suggested in the original papers, while only altering the replay buffer sampling method.
2.3 Correlating learning speed with catastrophic forgetting
@@
To simultaneously track learning speed and forgetting, we maintain the epoch-wise boolean classification matrix $M$ for each network (formally defined in §2.2). Throughout the training of both tasks, we record, for each epoch, whether the network correctly classified each example from the first task. The initial $E=100$ epochs represent classification during the first task, enabling us to observe the learning speed of the different examples, while the subsequent $E=100$ epochs allow us to monitor the forgetting of the first task during training of the second task.
In Figs. 2(b-c), we examine these epoch-wise classification matrices for both the train and test sets, respectively. To aid visualization, instead of plotting the epoch-wise classification matrix for each network, we aggregate together the networks from all runs using majority voting. The rows are sorted by the learning speed of the data samples. Notably, we observe a correlation between learning speed and catastrophic forgetting: examples learned more quickly during the first task tend to remain correctly classified throughout the second task, whereas slower-learned examples are more prone to immediate misclassification upon task transition, making them susceptible to forgetting. Essentially, slower learning speeds of examples correlate with a higher probability of forgetting by deep models.
@@
While the results presented here pertain to a simplified case, this phenomenon is robust across diverse datasets, architectures, and hyperparameters. For additional experiments, please refer to App. B.
2.4 How the replay buffer changes the catastrophic forgetting dynamics
(a) Small buffer, $4\%$
(b) Medium buffer, $12\%$
(c) Large buffer, $40\%$

### v2

# Predicting the Susceptibility of Examples to Catastrophic Forgetting

@@ Abstract
Catastrophic forgetting – the tendency of neural networks to forget previously learned data when learning new information – remains a central challenge in continual learning. In this work, we adopt a behavioral approach, observing a connection between learning speed and forgetting: examples learned more quickly are less prone to forgetting. Focusing on replay-based continual learning, we show that the composition of the replay buffer – specifically, whether it contains quickly or slowly learned examples – has a significant effect on forgetting. Motivated by this insight, we introduce Speed-Based Sampling (SBS), a simple yet general strategy that selects replay examples based on their learning speed. SBS integrates easily into existing buffer-based methods and improves performance across a wide range of competitive continual learning benchmarks, advancing state-of-the-art results. Our findings underscore the value of accounting for the forgetting dynamics when designing continual learning algorithms.
Machine Learning, ICML
@@ 1 Introduction
Recent years have seen a surge in the practical applications of deep learning, yet our mathematical and theoretical understanding of these models remains limited. While many advances are supported by some theoretical insights, the mechanisms behind key phenomena often remain elusive.
One such phenomenon is catastrophic forgetting (French, 1999; Kemker et al., 2018; McCloskey & Cohen, 1989), a core challenge in continual learning (CL) (see reviews: De Lange et al., 2021; Hadsell et al., 2020; Parisi et al., 2019). When models are exposed to new data, their performance on previously learned tasks often degrades. Despite extensive theoretical efforts, a full mathematical understanding of this behavior is still lacking.
In this work, we adopt a behavioral perspective, inspired by psychology and neuroscience, where internal mechanisms are often inaccessible. We train deep networks in various CL settings, observe their forgetting behavior, and infer insights from the resulting patterns.
Our key observation is a last-in-first-out forgetting pattern: examples learned later are more prone to forgetting, while earlier-learned ones are preserved. This aligns with the simplicity bias of neural networks (Shah et al., 2020; Szegedy et al., 2014), where simpler examples are learned first. As a result, simple examples are consistently remembered, while more complex ones are forgotten as new data arrives. This pattern holds across a wide range of architectures, datasets, and training configurations – including variations in learning rates, optimizers, schedulers, epochs, and regularization strategies (see §2, App. D). Fig. 1 visualizes remembered vs. forgotten examples in CIFAR-100.
We further explore how this effect changes as the model’s ability to remember examples improves. This improvement is achieved by either expanding the buffer size (§2.4), increasing the network size (App.C.3), or reducing dataset complexity (Fig.11). As memory improves, even slower-learned examples are remembered, leaving only the most complex ones susceptible to forgetting. Nonetheless, the core principle remains consistent: the slower an example is learned, the more likely it is to be forgotten.
Next, we examine how selective replay affects catastrophic forgetting. We rank examples based on their learning speed, and construct different buffer compositions by focusing on examples from a specific range of speeds. We find that selecting moderately fast-learning examples – the simplest and quickest to learn, which are still complex enough to benefit from replay – is particularly effective. As the model’s ability to remember improves, so does the optimal buffer focus shift progressively toward slower-to-learn examples.
Most competitive buffer-based methods rely on random sampling to fill the replay buffer, treating all examples from the old task equally. To demonstrate the potential of accounting for the susceptibility of examples to forgetting, we introduce a simple sampling strategy for replay buffers called speed-based sampling (SBS). SBS tracks the speed at which each example is learned during the old task training and prioritizes sampling the quickest-to-learn examples that are still prone to forgetting (details in §3.1). This approach is computationally and memory-efficient, requiring just one float per example computed during the forward pass, and integrates easily with any replay-based CL method. Despite its simplicity, SBS consistently improves performance across datasets and buffer sizes, demonstrating the practical value of our insights.
@@
Figure 1: Test examples that networks tend to either remember or forget. Networks were trained on 2 tasks from CIFAR-100. We show the examples that were learned quickest (left) and slowest (right). Slowest examples were often forgotten, while the quickest were rarely forgotten.
@@ 1.1 Our Contribution
  • Observing a connection between an example’s learning speed and its susceptibility to catastrophic forgetting.
  • We show that this observed relationship also persists across CL settings, even as models become more capable of remembering.
  • Demonstrating that this phenomenon is robust across diverse datasets and hyperparameters.
  • Showing that restricting the replay buffer to examples of specific learning speeds significantly affects forgetting across different continual learning scenarios.
  • Proposing a practical application of the observed phenomena, by introducing a novel sampling strategy for replay buffers, called speed-based sampling (SBS). SBS integrates with existing continual learning algorithms, significantly improving their performance across a wide range of methods and advancing the state of the art.
@@ 1.2 Related Work
Catastrophic forgetting. Most existing work on catastrophic forgetting focuses on mitigation (Kirkpatrick et al., 2017; Lee et al., 2017; Li et al., 2019; Ritter et al., 2018; Serra et al., 2018; Verwimp et al., 2025), rather than its characterization. Some works have explored catastrophic forgetting from a model-centric perspective: Nguyen et al. (2020); Ramasesh et al. (2021) showed differences in forgetting patterns across model layers, while Mirzadeh et al. (2020); Pfülb & Gepperth (2019) demonstrated the impact of training hyper-parameters. Nguyen et al. (2019) identified task complexity as a factor influencing forgetting rates. In contrast, our approach is data-centric and behavioral – we link forgetting to the speed at which individual examples are learned
Simplicity Bias and Selective replay. Previous studies have shown that neural networks tend to learn simple patterns before more complex ones (Cao et al., 2021; Gissin et al., 2020; Gunasekar et al., 2018; Heckel & Soltanolkotabi, 2020; Hu et al., 2020; Jin & Montúfar, 2023; Kalimeris et al., 2019; Pérez et al., 2019; Soudry et al., 2018; Ulyanov et al., 2018) – a phenomenon known as simplicity bias (Dingle et al., 2018; Shah et al., 2020). Our work extends this concept to catastrophic forgetting, revealing a reverse simplicity bias – complex examples are more likely to be forgotten than simple ones.
Our method leverages this insight by selectively replaying examples based on learning speed, indirectly controlling for complexity. Related paradigms such as curriculum learning (Bengio et al., 2009; Hacohen & Weinshall, 2019) and self-paced learning (Kumar et al., 2010), structure training around example complexity, demonstrating that such structure can improve generalization. However, unlike these approaches, we operate in a continual learning setting. We do not modify the original data or training sequence. Instead, we influence learning solely through selective replay, using example complexity as the criterion and observing its effect on catastrophic forgetting.
Our work also aligns with insights from semi-supervised learning, transfer learning, and active learning, where the complexity of the limited labeled data significantly affects algorithm design and performance (Hacohen et al., 2022; Yehuda et al., 2022; Hacohen & Weinshall, 2023a, b; Fluss et al., 2023; Chen et al., 2024). In all these settings, understanding and leveraging the distribution of example complexity can be key to effective learning.
Forgetting Dynamics. Several studies investigate forgetting dynamics during training. Maini et al. (2022); Toneva et al. (2019) show that, in non-continual learning settings, certain examples are inherently less prone to forgetting during training. Millunzi et al. (2023) highlighted distinct forgetting behaviors for noisy versus clean labels, proposing adjustments to the rehearsal process to account for these differences. Our findings build on this body of work by further identifying which examples are most susceptible to forgetting, linking catastrophic forgetting to simplicity bias. We propose a method to proactively identify examples susceptible to forgetting and demonstrate how these insights can be applied to optimize replay buffer sampling across various CL methods.
Replay buffer sampling functions. Replay-based methods mitigate forgetting by storing past examples in a fixed-size buffer. While numerous sampling strategies have been proposed (Aljundi et al., 2019; Benkő, 2024; Buzzega et al., 2021; Rebuffi et al., 2017; Tiwari et al., 2022; Wiewel & Yang, 2021), many are limited to specific conditions or methods. As a result, uniform sampling remains common (Buzzega et al., 2020; Guo et al., 2020; Kirkpatrick et al., 2017; Lopez-Paz & Ranzato, 2017; Prabhu et al., 2020; Ramesh & Chaudhari, 2022; Rolnick et al., 2019). In contrast, our proposed SBS offers a lightweight, general approach that consistently enhances replay-based methods across tasks, datasets, and buffer sizes.
2 Catastrophic Forgetting Vs. Learning Speed
@@
In this section, we observe the behavior of neural models when they exhibit catastrophic forgetting under different settings, drawing a connection between the speed at which a model learns examples, and its likelihood to forget them.
@@ 2.1 Definitions
We examine the CL classification setting, where the training dataset $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ consists of examples $x\in\mathcal{X}$ and their corresponding labels $y\in\mathcal{Y}$. This dataset is divided into $T>1$ tasks, with each task $t\in[T]$ represented as $\mathcal{D}_{t}=\{\mathcal{X}_{t},\mathcal{Y}_{t}\}$, and $\mathcal{D}=\bigcup_{t=1}^{T}{D_{t}}$. There is no overlap between data of different tasks. A deep model $f:\mathcal{X}\rightarrow\mathcal{Y}$ is sequentially trained on each task $\mathcal{D}_{t}$ from $t=1$ to $t=T$, with $E\in\mathbb{N}$ epochs per task. Each task also has a separate test set from the same distribution. During training on a task $\mathcal{D}_{t}$, the model $f$ is provided with a replay buffer $B$ of fixed-memory size, containing examples from previous tasks ${1,...,t-1}$. Notably, $|B|\ll|\mathcal{D}|$, meaning only a fraction of the data can be stored in the replay buffer. Note that the buffer may be empty ($B=\emptyset$), representing training $f$ without a replay buffer.
Learning speed. Neural networks demonstrate a trend in learning examples, where certain examples are consistently learned before others across various deep neural models (Baldock et al., 2021; Choshen et al., 2022; Hacohen et al., 2020). This orderliness is often connected to the simplicity bias phenomenon (Shah et al., 2020; Hacohen & Weinshall, 2022), as examples learned earlier are generally simpler than those learned later. Here, we measure the speed at which an example is learned, which also serves as a proxy for the example’s complexity. Formally, given a model $f$ and an example $(x,y)$ from either the training or test data, we define the learning speed of the example $(x,y)$ as:
@@
Here, $f_{t}^{e}$ denotes the intermediate model $f$ after training for $e\in[E]$ epochs on task $\mathcal{D}_{t}$. This definition is based on the accessibility score introduced by Hacohen et al. (2020), which we adapt for computation using a single model instead of an ensemble, facilitating efficient evaluation throughout the training of $f$. Intuitively, this score correlates with how quickly the model learns an example: if an example is correctly classified from the early stages of learning, it will sustain $f_{t}^{e}(x)=y$ for more epochs, resulting in a higher learning speed. A discussion of alternative metrics to learning speed can be found in App. G. A detailed explanation of the relationship between the example’s learning speed and this score is provided in Hacohen et al. (2020), and thus is not repeated here.
Computing learning speed. To calculate the learning speed of each example, we maintain a boolean matrix, called the epoch-wise classification matrix, $M\in\left\{0,1\right\}^{E\times|\mathcal{D}_{t}|}$ throughout the training on task $\mathcal{D}_{t}$, indicating whether the model correctly classified each example during learning. Specifically, for example $(x_{i},y_{i})\in\mathcal{D}_{t}$, we have $M_{e,i}=\mathds{1}[f_{t}^{e}(x_{i})=y_{i}]$. An example’s learning speed is the average of the $M$ across epochs. This matrix can be computed during the forward pass of the network, incurring minimal computation overhead. Furthermore, memory usage can be reduced by directly computing the matrix’s mean during training, maintaining a single float for each example.
@@ 2.2 Preliminaries
Class Incremental vs. Task Incremental Learning. Although similar, task incremental learning differs (TIL) from class incremental learning (CIL) by providing the task identity for each example during testing, allowing the use of the specific classifier for the classes within the task. We conducted experiments for both TIL and CIL. Our qualitative results across all experiments were consistent for both TIL and CIL. Therefore, to avoid redundancy, we present in the main paper results only for TIL setting, and repeat all the figures and experiments for CIL in App. B.
Datasets. We investigated various image continual learning classification tasks using split versions of several image datasets, including CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and TinyImageNet (Le & Yang, 2015). The data is split into $T$ tasks by partitioning the classes into $T$ equal-sized subsets. This partitioning is denoted as dataset-T. For example, splitting CIFAR-10 into 5 classes is denoted as CIFAR-10-5, comprising 5 tasks, each with 2 distinct classes. Unless otherwise specified, classes are divided into tasks according to the original order they appeared in the dataset (often alphabetically).
For detailed implementation details, including the architectures and hyper-parameters used, please see App. A.
2.3 Learning Speed vs. Catastrophic Forgetting
@@
To simultaneously track learning speed and forgetting, we maintain the epoch-wise boolean classification matrix $M$ for each network (formally defined in §2.2). Throughout the training of both tasks, we record, for each epoch, whether the network correctly classified each example from the first task. The initial $100$ epochs represent classification during the first task, enabling us to observe the learning speed of the different examples, while the subsequent $100$ epochs allow us to monitor the forgetting of the first task during the training of the second task.
In Figs. 2(b-c), we examine the epoch-wise classification matrices for the train and test sets, respectively. To aid visualization, instead of plotting the epoch-wise classification matrix for each network, we aggregate the networks from all runs using majority voting. The examples in the matrix are sorted by their learning speed. We observe a correlation between learning speed and catastrophic forgetting: examples learned more quickly during the first task tend to remain correctly classified throughout the second task, whereas slower-learned examples are prone to immediate misclassification upon task transition, making them susceptible to forgetting. Essentially, slower learning speeds of examples correlate with a higher probability of forgetting by deep models. Additional classification matrices, for different datasets and different amounts of tasks, can be found in App. C and Fig. 15.
@@
While the results presented here pertain to a simplified case, this phenomenon is robust across diverse datasets, architectures, and hyperparameters. For additional experiments, please refer to App. C, D.
2.4 Adding a Replay Buffer of Different Sizes
(a) Small buffer
(b) Medium buffer
(c) Large buffer